What would be an exact and approximate solution to 3 decimal places for x^2-5x+3=0
Answer by richwmiller(17219)

x^2-5x+3=0
complete the square.
(-5/2)^2=25/4
x^2-5x+25/4=25/4-12/4
x^2-5x+25/4=13/4
(x-5/2)^2=13/4
x = 1/2(5-sqrt(13))
x = 1/2(5+sqrt(13))
and approximate
x=0.697224=.697
x=4.30278=4.303